Anne borrows $1000 and agrees to pay $100 plus 1% interest if 1% on the unpaid balance each month. Find the payments for the first 6 months and the remaining debt at the end of this period.
Answer by edjones(8007)

100+1000*.01=110
100+900*.01=109
100+800*.01=108
+107+106+105
$400 balance at the end of 6 months.
.
Ed